OpenStudy (anonymous):

Consider the set of ordered pairs {(8, -1), (20, -10), (-12, 14)}. If they satisfy a linear relationship, what is the value of x when y = 60?

OpenStudy (phi):

If they satisfy a linear relationship means if they lie on the same line you have 3 points. You only need 2 of them to find the equation of a line through them

OpenStudy (phi):

you could try to find the equation y = m x + b where m is the slope, and b is the y-intercept or you could use the formula y - y0 = m( x - x0) where (x0,y0) is a point on the line, and m is the slope.

OpenStudy (amistre64):

a thought i have is to move this to the origin: zero out one of the points and move the rest by the same amount { (8, -1), (20, -10), (-12, 14) } ( x , 60) -8 +1 -8 +1 -8 +1 -8 +1 ----------------------------- ------- 0 0 12 -9 -20 15 x-8, 61 -9/12 -15/20 61/(x-8) they should all be equal -3/4 -3/4 proportions \[\frac{-3}{4}=\frac{61}{x-8}\] but tis is just one way to look at it

OpenStudy (anonymous):

wait so are you saying that the x value would have to be -3/4??